1000 raffle tickets are sold at $3 each. One grand prize of $100 and two consolation prizes of $75 will be awarded. Find Jake’s expectation if he purchases one ticket.
A) -$2.75 B) -$3.00 C) $0.24 D) $2.75 E) $3.00

Answer by feliz1965(151)

This is what I came up with:
250/x = 1000/1
Let x = what must be subtracted from the prize of one ticket
x(1000) = 250
x = 250/1000
x = 1/4 = 0.25
Then 3.00 - 0.25 = 2.75
If this is not correct, write back.